Given a mortgage of a $100000 at 7% for 15 years, prepare a one month amortization schedule. What is the balance after the 1st month's payment?
Answer #1

As per given information:

PrincipL (p)= 100000

R = 0.07/12

N = 15 X 12 = 180

Monthly installment = [P x R x (1+R)^N]/[(1+R)^N-1]

= [100000 X 0.07/12 x (1+0.07/12)^180] / [(1+0.07/12)^180-1]

= $898.83

Month Total Mortgage Payment Payment on Interest Payment on Principal Principal Balance
1 $898.83 100000*0.07/12 = $583.33 $898.83 - 583.33= $315.49 $100,000.00

Balance after the 1st month's payment = $100,000.00 - $315.49 = $996684.51